An alternating pd 220V amplitude and frequency 50Hz is applied to a series circuit containing R=400 ohms, L=0.35H and C= 18 micro farad (a) what is Z? (b) what is the current amplitude ,is it lagging the supply pd (c) what is the power factor?



1
Expert's answer
2021-02-15T17:40:50-0500

Given,

Potential difference (V)=220V

"(f)=50Hz"

"R=400 \Omega"

"L=0.35H"

"C= 18 \mu F"

"z=\sqrt{R^2+(X_L-X_C)^2}"


"=\sqrt{400^2+({2\pi fL-\frac{1}{2\pi f c})^2}}"


"=\sqrt{16\times 10^4+(2\pi \times 50\times 0.35-\frac{1}{2\pi \times 50\times 18\times 10^{-6}})^2}"


"=\sqrt{16\times 10^4+(35\pi-176.92})^2"


"=\sqrt{16\times10^4+(109.9-176.9)^2}"

"=\sqrt{16\times^4+4489}"

"=\sqrt{160000+4489}\Omega"

"=" "405.6\Omega"


"i=\frac{V_o}{Z}A"


"=\frac{220}{405.6}A"


"=0.54A"

Power factor "\cos(\phi) = \frac{R}{Z}"

"\cos\phi=\frac{400}{405.6}"

"=0.98"